Suppose that you’ve obtained the following information about Park Avenue Parking, Inc. Next year’s free cash flow to the firm is expected to be $7,500,000. Free cash flows are expected to grow at an annual rate of 2%. The firm’s equity beta is estimated to be 1.5, and the firm’s debt beta is estimated to be 0.3. The risk-free rate is 3% and the market risk premium is 6%. The market value of the firm’s debt is estimated to be $10,000,000, and the firm’s tax rate is 40%. Use iterative valuation to derive the current value of Park Avenue Parking.
Answer #1

Answer

Value of firm= Free cashflow(1+g)/(RE-g)

g= growth rate,=2%

free cash flow=7500000

RE= Required rate of return= RF+ Beta× market risk premium

=3+1.8×6=13.8

(RF= risk free rate of return=3%

market risk premium= 6%

Beta=equity beta+ debt beta=1.5+0.3= 1.8)

Now= value of firm=Free cashflow(1+g)/(RE-g)=7500000(1+0.02)/(13.8-2)=648305 answer